# Can Reasoning LLMs Enhance Clinical Document Classification?11footnotemark: 1

### v1

@@ Abstract
Methods: This study used the MIMIC IV dataset, consisting of 3,000 balanced discharge summaries 150 positive and 150 negative samples for each of the top 10 ICD-10 codes. The summaries were tokenized using cTAKES, which converted clinical narratives into structured SNOMED codes, capturing contextual details such as affirmation or negation. Eight LLMs, including four reasoning (Qwen QWQ, Deepseek Reasoner, GPT o3 Mini, Gemini 2.0 Flash Thinking) and four non-reasoning models (Llama 3.3, GPT 4o Mini, Gemini 2.0 Flash, Deepseek Chat), were evaluated over three experimental runs. Final predictions were determined using majority voting across the runs to assess accuracy, F1 score, and consistency.
@@
[label1]organization=College of Science and Engineering, James Cook University, addressline=, city=Townsville, postcode=4811, state=QLD, country=Australia
@@
[label2]organization=School of Computing, Macquarie University, addressline=, city=Sydney, postcode=2113, state=NSW, country=Australia
@@ 1 Introduction
Clinical document classification faces significant challenges due to the complex nature of medical narratives. The global clinical coding market size was valued at more than \$37 Billion USD in 2024 [1]. Clinical coding process faces significant challenges due to the complex, vary nature of medical narratives. These documents use heavy domain specific terminologies, abbreviations, and unique writing styles that vary by institution and practitioner. Moreover, in additional to healthcare strict privacy regulations, the limited availability of large, high quality annotated datasets slows down the development of robust classification models. The frequent imbalance in class distributions and the inherent ambiguity in clinical language further complicate model training and evaluation. To address these issues, specialized Natural Language Process (NLP) frameworks such as the clinical Text Analysis and Knowledge Extraction System (cTAKES) have been developed to better handle the unique characteristics of clinical texts [2].
@@
Clinical document classification is one of the processes that healthcare industry relays on to convert unstructured medical documents to International Classification of Diseases, 10th Revision (ICD-10) diagnoses [3]. ICD-10 coding has become a important process of clinical practice by providing a standardized language for documenting diagnoses and procedures. This systematic classification facilitates efficient communication among healthcare providers, ensuring that patient information is accurately recorded and easily shared across different institutions and systems [4, 5]. The consistent application of ICD-10 codes not only enhances clinical documentation and billing processes but also plays a pivotal role in epidemiological research, quality assurance, and public health reporting [6]. By enabling comprehensive tracking of disease trends and treatment outcomes, ICD-10 coding informs health policy and resource allocation, ultimately contributing to improved patient care and safety [7, 8].
@@
Large language models (LLMs) have rapidly emerged as a transformative technology in healthcare, driven by their advanced natural language understanding and reasoning capabilities [9]. These models, trained on extensive and diverse corpora that include medical literature and clinical narratives, can parse complex clinical language, infer relationships among diverse medical concepts, and generate contextually nuanced insights [10, 11]. Such reasoning abilities are particularly beneficial for automating tasks like clinical document classification and diagnosis coding, where they can support decision making and improve the efficiency and accuracy of healthcare delivery. Recent developments in LLM architectures, including techniques like chain-of-thought prompting, have further enhanced their capacity to handle the intricacies of clinical data, marking a significant shift from traditional NLP approaches [12, 13].
@@
Comparing reasoning versus non-reasoning large language models addresses a crucial research gap in applying these systems to complex clinical tasks. While models enhanced with explicit reasoning have shown promise in capturing intricate medical relationships and nuances in clinical narratives [14, 15], their benefits over standard non-reasoning models have not been systematically quantified in clinical document classification. This comparison is essential to determine whether the additional computational complexity and resource demands of reasoning based approaches translate into significantly improved diagnostic coding accuracy and consistency, ultimately informing more effective and interpretable model deployment in healthcare settings.
@@
This study addresses several critical research questions that bridge the gap between reasoning based and non-reasoning LLMs in clinical document classification. First, it investigates how reasoning based LLMs compare with their non-reasoning counterparts in classifying ICD-10 coded discharge summaries, exploring whether the enhanced reasoning abilities translate to improved handling of clinical narratives. Second, the research evaluates the overall classification accuracy of LLMs on tokenized discharge summaries, providing quantitative insights into their performance on a challenging and balanced dataset. Third, the study examines the consistency of classification results across multiple experimental runs for each model, thereby assessing the robustness and reliability of these approaches. Collectively, these research questions aim to illuminate the strengths and limitations of current LLM methodologies in healthcare, offering valuable contributions toward the optimal deployment of AI in clinical settings.
@@ 2.1 Clinical Text Classification
Recent advances in clinical text classification have progressively shifted from traditional machine learning techniques to more nuanced deep learning methods that better capture the complexity of clinical language. For instance, the work by Haoran Shi et al. [16] introduces a hierarchical deep learning model with an attention mechanism specifically designed for the automated assignment of ICD diagnostic codes from written diagnosis descriptions. Their approach leverages character aware neural language models to generate detailed hidden representations of both clinical narratives and ICD codes, effectively addressing the inherent mismatch between the number of descriptions and their corresponding codes. The significant performance gains, marked by an F1 score of 53% and an AUC of 90%, underscore the potential of deep learning frameworks over conventional methods that often rely on manual feature engineering. This research exemplifies how integrating attention mechanisms within deep learning architectures can enhance the accuracy and reliability of clinical text classification, setting a new benchmark for subsequent studies in the field.
2.2 ICD Coding & MIMIC Dataset
@@ 2.3 Tokenization & Preprocessing in Clinical NLP
In clinical NLP, the preprocessing of unstructured medical texts is essential for ensuring that downstream models can effectively interpret clinical narratives. In his work on classifying medical notes into standard disease codes [18], Karmakar emphasizes a systematic approach to note preprocessing that includes standard steps such as lowercasing, removal of special characters, and tokenization. Tools like cTAKES are particularly relevant in this context because they provide domain specific processing capabilities tailored to the nuances of clinical language. cTAKES not only streamlines tokenization and sentence boundary detection but also offers robust named entity recognition that can identify medical terms, abbreviations, and concepts despite variations in clinical documentation. This targeted preprocessing is crucial for managing the inherent variability of discharge summaries and ultimately contributes to the effectiveness of models, such as CNNs and Long Short Term Memory models (LSTMs), when classifying complex clinical data into ICD codes.
@@ 2.4 Large Language Models for Clinical Coding
Rumeng Li et al.’s work [19] provides a compelling comparison between reasoning and non-reasoning approaches within LLMs for ICD coding. Their study demonstrates that while traditional non-reasoning models and straightforward prompting methods often yield high recall, they tend to struggle with precision largely due to issues such as hallucination and the inherent complexity of mapping clinical narratives to a vast and skewed label space. In contrast, reasoning based techniques, notably those leveraging chain-of-thought prompting and self consistency, introduce multi-step reasoning processes that enhance the model’s interpretability and robustness. However, even these reasoning models fall short when used in isolation, as they can still produce incomplete or inaccurate results. To address these challenges, Li et al. propose a multi-agent system that simulates the real world ICD coding workflow, effectively combining the strengths of reasoning based methods with domain specific strategies such as integrating the Simple Object Access Protocol (SOAP) structure to achieve superior performance, particularly in handling rare codes. This comparison underscores the necessity of incorporating structured reasoning and interactive agent dynamics in LLMs to fully harness their potential for complex, domain specific tasks like ICD coding.
Mustafa et al.’s work [20] is closely aligned with the ongoing exploration of LLMs for clinical coding. By assessing ChatGPT 3.5 and 4 in comparison to traditional machine learning and SNOMED mapping methods, thier study addresses the challenge of accurately classifying ICD-10 codes in complex medical records. While human coders still demonstrated superior accuracy, ChatGPT 4’s ability to match the median human performance and exhibit improved consistency suggests that advanced LLMs can play a valuable role in clinical coding. This research highlights the potential of LLMs to enhance coding accuracy and efficiency while also emphasizing the variability in performance across different ICD-10 codes, key considerations for integrating AI into healthcare documentation workflows.
@@ 3.1 Dataset Preparation
The analysis incorporated both primary and secondary diagnoses, forming a set of the top 10 ICD-10 codes. This dual list approach was selected to capture a more comprehensive spectrum of clinically relevant conditions. By combining both primary and secondary diagnoses, the study benefits from a richer diagnostic landscape, which is essential for robust evaluation of classification performance across diverse clinical scenarios. To ensure class balance within the dataset, for each of these 10 ICD-10 codes, 150 discharge summaries confirmed as positive (i.e., diagnosed with the respective ICD-10 code) were extracted. An equal number of 150 negative discharge summaries, not associated with the given ICD-10 code, were randomly sampled. This approach yielded a total of 300 discharge summaries per ICD-10 code, resulting in an aggregate dataset of 3,000 discharge summaries.
@@
Table 1: Top 5 primary diagnoses cases and top 5 case count and their respective total cases in ICD-10.
@@ 3.2 Preprocessing
cTAKES was utilized to convert unstructured discharge summaries into structured, machine interpretable data by transforming the clinical narratives into lists of standardized SNOMED codes. These codes represent a comprehensive array of clinical entities such as diseases, lab tests, procedures, medications, symptoms, anatomy, and events. During the tokenization process, cTAKES not only extracted these terms but also determined their contextual status categorizing each as either affirmed or negated. After processing, the extracted SNOMED items were aggregated into two separate sections based on their contextual designation, and the frequency of each term’s occurrence was recorded. This comprehensive approach provided a robust dataset of tokenized discharge summaries, enriched with detailed frequency counts and contextual annotations, which served as the foundation for subsequent classification experiments using large language models.
@@ 3.3 Experimental Setup
For the experimental procedure, each of the eight models was tasked with classifying a dataset of 3,000 tokenized discharge summaries that had been preprocessed using cTAKES. To ensure the robustness and repeatability of the results, each model was run three times, with each run independently processing the entire dataset. In every experimental iteration, the models generated classifications indicating whether each discharge summary belonged to its corresponding ICD-10 group. The final classification outcome for each summary was determined using a majority vote across the three repetitions, meaning that the label assigned by at least two out of three runs was taken as the definitive result. This methodological approach not only ensured consistency in the classification outcomes but also provided a reliable measure of each model’s performance in distinguishing between the ICD-10 groups. Figure 1 illustrates this process.
@@ 3.4 Evaluation Metrics
Lastly, we assessed how consistently each model classified the same discharge summaries across its three runs. Specifically, for each discharge summary, we recorded whether the model’s prediction (positive or negative) remained the same in all three trials. The proportion of summaries that received identical classifications in each run was then computed as an indicator of consistency. This measure offers insight into the reliability of the model’s decision making process, revealing whether fluctuations in output are minimal (high consistency) or more pronounced (low consistency) over repeated evaluations.
Data and scripts are available on GitHub: https://github.com/asmgx/LLMs.
@@ 4 Results
4.1 Performance Metrics
Among the eight LLM models evaluated, Gemini 2.0 Flash Thinking, a reasoning model, achieved the highest overall accuracy at 75% and the highest F1 score at 76%. In contrast,GPT 4o Mini, a non-reasoning model, had the lowest performance among all models, with an accuracy of 64% and an F1 score of 47%. Among the non-reasoning models, Gemini 2.0 Flash recorded the highest accuracy at 72% and the highest F1 score at 71%, while Qwen QWQ had the lowest accuracy and F1 score among reasoning models, with 65% accuracy and 55% F1 score. Table 3 presents the detailed accuracy and F1 score results for each model.
Table 3: Large Languages Models used.
@@ 4.1 Performance Metrics
Overall, the reasoning models outperformed the non-reasoning models in terms of average accuracy and F1 scores. The reasoning models achieved an average accuracy of 71% compared to 68% for the non-reasoning models. Similarly, the reasoning models recorded an average macro F1 score of 67%, whereas the non-reasoning models achieved a lower average macro F1 score of 60%, Figure 2 shows the results. These results suggest that, in general, the reasoning models provided slightly better classification performance than the non-reasoning models.
Figure 2: Measures by Reasoning Type.
@@ 4.2 Consistency Analysis
The consistency of model performance across the three experiments reveals a varied level of reliability among the different LLMs. Models such as Deepseek GPT o3 Mini at 95%, Chat at 95%, and Gemini 2 Flash at 91% demonstrate high consistency, with performance largely stable across all three trials. This suggests that these models are robust in classifying discharge summaries, maintaining similar results across multiple runs. On the other hand, models like Qwen QWQ and Gemini 2 Flash Thinking, with a consistency rate of 78%, show significantly lower consistency, indicating greater variability in their performance across the experiments. This variability could potentially indicate issues with stability or sensitivity to different input data or model configurations. Other models such as Deepseek Reasoner at 83%, Llama 3.3 at 88%, and GPT 4o Mini at 89% fall in between, suggesting a moderate level of consistency. Overall, the high consistency observed in certain models like Deepseek Chat and GPT o3 Mini suggests they are more reliable for repeated use in clinical document classification tasks. These results are shown in Table 3. An analysis of the average consistency based on model type, reveals that non-reasoning models exhibit a higher average consistency of 91% compared to reasoning models, which show an average consistency of 84%. Figure 2 illustrates these findings in detail.
4.3 Performance Analysis of LLMs on ICD-10 Code Classification
Analyzing the performance of LLMs on ICD-10 codes using radar charts displayed in Figures 3 to 12, we can classify these codes into three ICD-10 categories. The first category consists of ICD-10 codes where all LLMs follow a similar trend in accuracy and F1 score. The second category includes codes where there is a significant difference between accuracy and F1 score, indicating inconsistencies in positive case identification. The third category consists of ICD-10 codes where most LLMs fail to identify positive discharge summaries, leading to highly skewed results.
In the first category, ICD-10 codes such as A41, I10, I21, and I25 exhibit a close match between accuracy and F1 score across all LLMs. Among these, reasoning LLMs such as Deepseek Reasoner, Gemini 2 Flash Thinking, and GPT 3o Mini performed exceptionally well, achieving over 80% accuracy and F1 score for all four codes. Particularly for I25, almost all LLMs demonstrated strong performance, with seven out of eight models exceeding 90% accuracy and F1 score. The only exception was Llama 3.3, which had slightly lower results of 87% accuracy and 89% F1 score.
The second category includes ICD-10 codes E78 and I13, where certain LLMs showed a considerable gap between accuracy and F1 score. For example, GPT 4o Mini had a significant disparity, with an accuracy of 56% but an F1 score of only 23% for E78, and 58% accuracy with an F1 score of 30% for I13. Other models, such as Deepseek Chat and Qwen QWQ, also exhibited accuracy and F1 score variances, though not as extreme. This discrepancy often occurs when the number of true negative records is significantly higher than the number of true positive records. In these cases, models tend to classify most discharge summaries as negative, even when some should be identified as positive, leading to misleadingly high accuracy scores.
The third category consists of ICD-10 codes Y92, Z51, Z79, and Z87. Here, only Deepseek Reasoner, Gemini 2 Flash, and Gemini 2 Flash Thinking managed to identify a small number of positive discharge summaries, while all other models failed almost entirely. In particular, GPT 4o Mini, Deepseek Chat, and GPT 3o Mini classified all summaries as negative for Z87. Similarly, for Y92, which specifies the location of an event rather than a diagnosis, most models failed to distinguish between relevant and irrelevant cases [21]. This suggests that without proper training, LLMs may misclassify notes based on contextual information such as place of occurrence or patient history rather than actual medical conditions. Tables from 4 to 13 show these results in details.

### v2

@@ Abstract
Methods: This study used a balanced subset of the MIMIC-IV dataset, comprising 3,000 discharge summaries including 150 positive and 150 negative samples for each of the top 10 ICD-10 codes. The summaries were tokenized using cTAKES, which converted clinical narratives into structured SNOMED codes, capturing contextual details such as affirmation or negation. Eight LLMs, including four reasoning (Qwen QWQ, Deepseek Reasoner, GPT o3 Mini, Gemini 2.0 Flash Thinking) and four non-reasoning models (Llama 3.3, GPT 4o Mini, Gemini 2.0 Flash, Deepseek Chat), were evaluated over three experimental runs. Final predictions were determined using majority voting across the runs to assess accuracy, F1 score, and consistency.
@@
[label1]organization=College of Science and Engineering, James Cook University, city=Townsville, postcode=4811, state=QLD, country=Australia
@@
[label2]organization=School of Computing, Macquarie University, city=Sydney, postcode=2113, state=NSW, country=Australia
@@ 1 Introduction
Clinical document classification faces significant challenges due to the complex nature of medical narratives. The global clinical coding market size was valued at more than \$37 Billion USD in 2024 [1]. The clinical coding process faces significant challenges due to the complex, varied nature of medical narratives. These documents use heavy domain specific terminologies, abbreviations, and unique writing styles that vary by institution and practitioner. Moreover, in addition to healthcare strict privacy regulations, the limited availability of large, high quality annotated datasets slows down the development of robust classification models. The frequent imbalance in class distributions and the inherent ambiguity in clinical language further complicate model training and evaluation. To address these issues, specialized Natural Language Processing (NLP) frameworks such as the clinical Text Analysis and Knowledge Extraction System (cTAKES) have been developed to better handle the unique characteristics of clinical texts [2].
@@
Clinical document classification is one of the processes that the healthcare industry relies on to convert unstructured medical documents to International Classification of Diseases, 10th Revision (ICD-10) diagnoses [3]. ICD-10 coding has become an important process of clinical practice by providing a standardized language for documenting diagnoses and procedures. This systematic classification facilitates efficient communication among healthcare providers, ensuring that patient information is accurately recorded and easily shared across different institutions and systems [4, 5]. The consistent application of ICD-10 codes not only enhances clinical documentation and billing processes but also plays a pivotal role in epidemiological research, quality assurance, and public health reporting [6]. By enabling comprehensive tracking of disease trends and treatment outcomes, ICD-10 coding informs health policy and resource allocation, ultimately contributing to improved patient care and safety [7, 8].
@@
Large language models (LLMs) have rapidly emerged as a transformative technology in healthcare, driven by their advanced natural language understanding and reasoning capabilities [9]. These models, trained on extensive and diverse corpora that include medical literature and clinical narratives, can parse complex clinical language, infer relationships among diverse medical concepts, and generate contextually nuanced insights [10, 11]. Such reasoning abilities are particularly beneficial for automating tasks like clinical document classification and diagnosis coding, where they can support decision-making and improve the efficiency and accuracy of healthcare delivery. Recent developments in LLM architectures, including techniques like chain-of-thought processing, have further enhanced their capacity to handle the intricacies of clinical data, marking a significant shift from traditional NLP approaches [12, 13].
@@
Comparing reasoning versus non-reasoning large language models addresses a crucial research gap in applying these systems to complex clinical tasks. While models enhanced with explicit reasoning have shown promise in capturing intricate medical relationships and nuances in clinical narratives [14, 15], their benefits over standard non-reasoning models have not been systematically quantified in clinical document classification. This comparison is essential to determine whether the additional computational complexity and resource demands of reasoning-based approaches translate into significantly improved diagnostic coding accuracy and consistency, ultimately informing more effective and interpretable model deployment in healthcare settings.
@@
This study addresses several critical research questions that bridge the gap between reasoning and non-reasoning LLMs in clinical document classification. First, it investigates how reasoning LLMs compare with their non-reasoning counterparts in classifying ICD-10 coded discharge summaries, exploring whether the enhanced reasoning abilities translate to improved handling of clinical narratives. Second, the research evaluates the overall classification accuracy of LLMs on tokenized discharge summaries, providing quantitative insights into their performance on a challenging and balanced dataset. Third, the study examines the consistency of classification results across multiple experimental runs for each model, thereby assessing the robustness and reliability of these approaches. Collectively, these research questions aim to illuminate the strengths and limitations of current LLM methodologies in clinical document classification, offering valuable contributions toward the optimal deployment of AI in clinical settings.
@@ 2.1 Clinical Text Classification
Recent advances in clinical text classification have progressively shifted from traditional machine learning techniques to more nuanced deep learning methods that better capture the complexity of clinical language. For instance, the work by Haoran Shi et al. [16] introduces a hierarchical deep learning model with an attention mechanism specifically designed for the automated assignment of ICD diagnostic codes from written diagnosis descriptions. Their approach leverages character-aware neural language models to generate detailed hidden representations of both clinical narratives and ICD codes, effectively addressing the inherent mismatch between the number of descriptions and their corresponding codes. The significant performance gains, marked by an F1 score of 53% and an AUC of 90%, underscore the potential of deep learning frameworks over conventional methods that often rely on manual feature engineering.
@@ 2.2 Tokenization & Preprocessing in Clinical NLP
In clinical NLP, the preprocessing of unstructured medical texts is essential for ensuring that downstream models can effectively interpret clinical narratives. In his work on classifying medical notes into standard disease codes [18], Karmakar emphasizes a systematic approach to note preprocessing that includes standard steps such as lowercasing, removal of special characters, and tokenization. Tools like cTAKES are particularly relevant in this context because they provide domain-specific processing capabilities tailored to the nuances of clinical language. cTAKES not only streamlines tokenization and sentence boundary detection but also offers robust named entity recognition that can identify medical terms, abbreviations, and concepts despite variations in clinical documentation. This targeted preprocessing is crucial for managing the inherent variability of discharge summaries and ultimately contributes to the effectiveness of models, such as CNNs and Long Short Term Memory models (LSTMs), when classifying complex clinical data into ICD codes.
@@ 2.3 Large Language Models for Clinical Coding
Li et al.’s work [19] provides a study exploring the use of multiple LLM agents collaborating to assign ICD codes to clinical notes. In this framework, distinct agents were tasked with reading clinical notes, generating relevant ICD codes, and then engaging in a discussion to reach a consensus. By structuring the interaction as a multi-agent dialogue, the authors aimed to simulate expert deliberation and improve coding accuracy. Their findings demonstrated that this collaborative LLM approach significantly enhanced performance, highlighting the potential of agent-based coordination in complex medical classification tasks..
Mustafa et al.’s work [20] is closely aligned with the ongoing exploration of LLMs for clinical coding. By assessing ChatGPT 3.5 and 4 compared to traditional machine learning and SNOMED mapping methods, their study investigates the challenge of accurately classifying ICD-10 codes in complex medical records. They demonstrate that, while human coders still achieve superior accuracy, ChatGPT 4’s ability to match the median human performance and exhibit improved consistency suggests that advanced LLMs can play a valuable role in clinical coding.
These studies highlight both the potential and challenges of using LLMs for clinical coding. While they show promising performance, issues such as hallucination and handling complex imbalanced code sets remain. Approaches that incorporate reasoning, domain knowledge, and structured workflows offer a path forward. In this research, we investigate if reasoning LLMs can enhance document classification.
@@ 3.1 Dataset Preparation
The analysis incorporated both primary and secondary diagnoses, forming a set of the top 10 ICD-10 codes. This dual list approach was selected to capture a more comprehensive spectrum of clinically relevant conditions. By combining both primary and secondary diagnoses, the study benefits from a richer diagnostic landscape, which is essential for robust evaluation of classification performance across diverse clinical scenarios. To ensure class balance within the dataset, for each of these 10 ICD-10 codes, 150 discharge summaries confirmed as positive (i.e., diagnosed with the respective ICD-10 code) were extracted. An equal number of 150 negative discharge summaries, not associated with the given ICD-10 code, were randomly sampled from the set of ICD codes, excluding the 10 selected codes. This approach yielded a total of 300 discharge summaries per ICD-10 code, resulting in an aggregate dataset of 3,000 discharge summaries.
@@
Table 1: Top 10 ICD codes used in this study. The top part of the table shows 5 codes with the highest number of primary diagnoses cases, while the bottom part shows the top 5 codes with the maximum total cases in the MIMIC IV dataset.
@@ 3.2 Preprocessing
cTAKES was utilized to convert unstructured discharge summaries into structured, machine-interpretable data by transforming the clinical narratives into lists of standardized SNOMED codes. These codes represent a comprehensive array of clinical entities such as diseases, lab tests, procedures, medications, symptoms, anatomy, and events. During the tokenization process, cTAKES not only extracted these terms but also determined their contextual status categorizing each as either affirmed or negated. This automated pipeline significantly reduced the length of the original documents by eliminating unnecessary text, which in turn improved downstream processing efficiency and cost. After processing, the extracted SNOMED items were aggregated into two separate sections based on their contextual designation, and the frequency of each term’s occurrence was recorded. This comprehensive approach provided a robust dataset of tokenized discharge summaries, enriched with detailed frequency counts and contextual annotations, which served as the foundation for subsequent classification experiments using large language models.
@@ 3.3 Experimental Setup
For the experimental procedure, each of the eight models was tasked with classifying a dataset of 3,000 tokenized discharge summaries that had been preprocessed using cTAKES. To ensure the robustness and repeatability of the results, each model was run three times, with each run independently processing the entire dataset. In every experimental iteration, the models generated classifications indicating whether each discharge summary belonged to its corresponding ICD-10 group. The prompt used for all models was:
“Discharge Summary: [Report Text] Does this summary contain the diagnosis associated with ICD-10 code [ICD10 Code] or any of its specific subcategories, Answer with Yes or No only?”
The final classification outcome for each summary was determined using a majority vote across the three repetitions, meaning that the label assigned by at least two out of three runs was taken as the definitive result. This methodological approach not only ensured consistency in the classification outcomes but also provided a reliable measure of each model’s performance in distinguishing between the ICD-10 groups. Figure 1 illustrates this process. In addition, in favour of reproducible research, our code and data are shared publicly on GitHub: https://github.com/asmgx/LLMs.
@@ 3.4 Evaluation Metrics
Lastly, we assessed how consistently each model classified the same discharge summaries across its three runs. Specifically, for each discharge summary, we recorded whether the model’s prediction (positive or negative) remained the same in all three trials. The proportion of summaries that received identical classifications in all runs was then computed as an indicator of consistency. This measure offers insight into the reliability of the model’s decision-making process, revealing whether fluctuations in output are minimal (high consistency) or more pronounced (low consistency) over repeated evaluations.
@@ 4 Results
4.1 LLMs Performance
Among the eight LLMs evaluated, Gemini 2.0 Flash Thinking, a reasoning model, achieved the highest overall accuracy at 75% and the highest F1 score at 76%. In contrast, GPT 4o Mini, a non-reasoning model, had the lowest performance among all models, with an accuracy of 64% and an F1 score of 47%. Among the non-reasoning models, Gemini 2.0 Flash recorded the highest accuracy at 72% and the highest F1 score at 71%, while Qwen QWQ had the lowest accuracy and F1 score among reasoning models, with 65% accuracy and 55% F1 score. Table 3 presents the detailed accuracy and F1 scores for each model.
Table 3: Average performance of various LLMs over 3 runs in classifying 3000 clinical discharge summaries from the top 10 ICD-10 codes within the MIMIC IV dataset.
@@ 4.1 LLMs Performance
Overall, the reasoning models outperformed the non-reasoning models in terms of average accuracy and F1 scores. The reasoning models achieved an average accuracy of 71% compared to 68% for the non-reasoning models. Similarly, the reasoning models recorded an average macro F1 score of 67%, whereas the non-reasoning models achieved a lower average macro F1 score of 60%. Figure 2 shows the average performance of the four state-of-the-art reasoning versus four non-reasoning LLMs in classifying clinical documents. These results suggest that, in general, the reasoning models provided slightly better classification performance than the non-reasoning models.
Figure 2: Average performance of 4 reasoning vs 4 non-reasoning LLMs in classifying clinical documents.
@@ 4.2 Consistency Analysis
The consistency of model performance across the three experiments reveals a varied level of reliability among the different LLMs. Results are shown in Table 3. Models such as GPT o3 Mini and Deepseek Chat at 95%, and Gemini 2 Flash at 91% demonstrate high consistency, with performance largely stable across all three trials. This suggests that these models are robust in classifying discharge summaries, maintaining similar results across multiple runs. On the other hand, models like Qwen QWQ and Gemini 2 Flash Thinking, with a consistency rate of 78%, show significantly lower consistency, indicating greater variability in their performance across the experiments. This variability could potentially indicate issues with stability or sensitivity to different input data or model configurations. Other models such as Deepseek Reasoner at 83%, Llama 3.3 at 88%, and GPT 4o Mini at 89% fall in between, suggesting a moderate level of consistency. Overall, the high consistency observed in certain models like Deepseek Chat and GPT o3 Mini suggests they are more reliable for repeated use in clinical document classification tasks. An analysis of the average consistency based on model reasoning, reveals that non-reasoning models exhibit a higher average consistency of 91% compared to reasoning models, which show an average consistency of 84%. Figure 2 illustrates these findings in detail.
4.3 Detailed Performance Analysis of Various LLMs
The performance analysis of the eight investigated LLMs in classifying ICD-10 codes is presented in two different ways. Tables 4 to 13 list the accuracy, F1 score, and consistency of all LLMs on each specific code. Additionally, Figures 3 to 12 demonstrate the F1 score of all LLMs vs their consistency, for each code. These results assist in classifying the codes into three categories.
The first category includes codes for which all models demonstrate strong performance in both consistency and F1 scores, as shown in Figures 3, 5, 7, and 8, exhibiting similar trends and levels of accuracy. This category comprises ICD-10 codes A41, I10, I21, and I25. For these codes, consistency ranges from 67% to 100%, while F1 scores range from 63% to 96%, as shown in Tables 4, 6, 8, and 9. Reasoning models, particularly GPT 3o Mini, stood out in terms of consistency, achieving the highest levels for 3 out of the 4 codes. In terms of F1 scores, both reasoning and non-reasoning models led, depending on the specific ICD code. Notably, GPT 4o Mini achieved the highest F1 score of 96%.
The second category includes codes where consistency remained high and stable across all models, but F1 scores varied significantly, as shown in Figures 4, 6, 9, and 10. This suggests that while the models consistently made similar predictions, their accuracy in identifying true positive cases varied widely. ICD-10 codes in this category include E78, I13, Y92, and Z51. Consistency scores ranged from 64% to 100%, while F1 scores showed a much wider spread—ranging from 23% to 79% for E78, 30% to 86% for I13, 0% to 20% for Y92, and 4% to 87% for Z51, as shown in Tables 5, 7, 10, and 11. GPT 4o Mini showed notably poor performance in this category, recording the lowest F1 scores for E78, I13, and Y92 at 23%, 30%, and 0%, respectively.
The third category includes codes Z79 and Z87, where both consistency and F1 scores varied greatly across models, indicating inconsistent performance in identifying positive discharge summaries, as shown in Figures as shown in Figures 11, 12. For Z79, consistency ranged from 33% to 100%, while F1 scores ranged from 0% to 57%. For Z87, consistency ranged from 36% to 100%, and F1 scores ranged from 0% to 62%, as detailed in Tables 12 and 13. GPT 3o Mini demonstrated 100% consistency for both codes but scored 0% on F1, indicating that it consistently misclassified all records as negative. Similarly, Deepseek Chat and GPT 4o Mini also failed to identify any positive cases for Z87, both recording an F1 score of 0%, although their consistency levels were slightly lower.